OpenStudy (anonymous):

NEED THIS DONE PLEASE HELP!!! ONE MEDAL Which is a solution of the equation? 2x + y = 5

OpenStudy (anonymous):

y = 5-2x

OpenStudy (anonymous):

This equation has infinite solutions. In order to solve it, we can assume x = constant = c Then y = 5 - 2x = 5 - 2c So solution set is (x,y) = (c, 5 - 2c), with c is a real number. Some examples in this set may be: (1, 3); (0, 5); etc.